OpenStudy (anonymous):

Write the equation of the line given the slope and a point on the line. Type your answer in the box without skipping any spaces. m = 7, point (-2, 0)

OpenStudy (anonymous):

use y-y1=m(x-x1) so it will be y-0=7(x+2) the equation will then be y=7x+14